(3x+3y-1)dx+(x+y+1)dy=0 solve the separable equation using linear substitution


1
Expert's answer
2021-11-26T00:07:12-0500

"Solution: Given ~~
\\(3x+3y-1)dx+(x+y+1)dy=0\\Now,Substitute~\\ u=x+y~ \implies du=dx+dy ~\\y=u-x~\implies dy=du-dx
\\Put~ in~ given ~ordinary ~~differential~ ~equation
\\(3(u-x)+3x-1)dx+(u+1)(du-dx)=0 \\\implies (3u-3x+3x-1)dx+udu-udx-du-dx=0
\\Simplifying,we~ get
\\(u+1)du=(2-2u)dx
\\Divide ~by ~2-2u
\\\frac{(-u-1)}{2(u-1)}du=dx
\\Which ~is~ separable~ differential~ equation~form~ M(x)du=N(x)dx
\\ \therefore Integrating~both~sides~
\\\int \frac{(-u-1)}{2(u-1)}du=\int dx
\\\therefore \frac{-u}{2}-ln ~(u-1)=x+c
\\back ~substitution~u=x+y
\\-ln (x+y-1)-\frac{x+y}{2}=x+c
\\Simplify, ~we ~get
\\-ln (x+y-1)-\frac{y}{2}=frac{3x}{2}+c,x=1-y"